OpenStudy (anonymous):

If y=m^4=n^3 and y is greater than 1, then mn=

OpenStudy (anonymous):

y^(1/2) y^(1/7) y^(7/12) y^7 y^12

OpenStudy (zzr0ck3r):

y=m^4 imples m=y^(1/4) y=n^3 implies n=y^(1/3) n*m = y^(1/3)*y^(1/4)=y^(1/3+1/4) = y^(7/12)

OpenStudy (zzr0ck3r):

we use the fact that y is positive to take the fourth root